OpenStudy (anonymous):

what is the derivative of -2(ln(3^4x))?

OpenStudy (anonymous):

factor out your constant: -2((d/dx)ln(3^(4x))) use the chain rule: ((dln(u))/du) (du/dx) where u = 3^(4x) use the chain rule again: ((d/dx) 3^4x (d3^u)/du (du/dx) where u = 4x factor out your constants: -2ln(3)(4(d/dx (x))) derivative of x = 1 d/dx(-2ln(3^(4x))) = -8ln(3)

OpenStudy (anonymous):

thanks!! that is what I got, but my answer key says otherwise.

OpenStudy (anonymous):

yeah, you may simplify it further, or may have simplified differently during an intermediate step